OpenStudy (swag):

100 plus the product of a number and -2 equals 50. What is the number?

OpenStudy (klimenkov):

\(100+(-2)\cdot x=50\) \(x\) is the number.

OpenStudy (klimenkov):

\[\Large{x=5\,^{\int \limits_0^{\pi}\sin xdx}}\]

OpenStudy (swag):

So it would be 25?

OpenStudy (klimenkov):

Yes.

OpenStudy (swag):

Thank You